# Deep Spherical Superpixels

### v1

@@ 1 Introduction
The vast majority of computer vision methods are tailored for standard RGB images, i.e., captured with a standard 90° field of view (FoV). However, acquisition devices with wider FoV have become more and more popular in the recent years. In particular, omnidirectional images with a 360°$\times$180° FoV are very interesting to capture the entire environment of a scene. Over the literature, such imagery may be equally referred as omnidirectional, spherical, 360°, or even panoramic. Naturally, such acquisition introduces distortions when projecting the capture on a planar 2D image. Nevertheless, many dedicated methods have been successfully applied on these images, for example for scene reconstruction [21], semantic segmentation [27] for autonomous driving, or in the context of mixed or virtual reality [18].
@@ 2.1 Superpixel Sampling Network
The SSN model takes as input images of size $N=h\times w$, represented with 5 channels corresponding to Lab color features (3 channels) and $xy$ pixel coordinates (2 channels). The goal of the model is to learn deep features that are more suitable to perform a differentiable clustering into superpixels. To achieve this, the SSN model uses a CNN composed of three blocks, each with two convolutional layers, batch normalization and ReLU activation, with a max pooling layer applied after each block. For the output, feature maps of each block are upsampled to the original image size (if necessary, for the second and third blocks) and concatenated. The original Lab and xy features are also concatenated into the output feature maps, resulting in $D$-dimensional pixel features (i.e., 5 channels from input features and $D-5$ learned deep features). In practice, the SSN model used $D=20$ in their experiments. For more details about this architecture, the reader can refer to [12].
@@ 2.3.1 Features and superpixels initialization
As input for the CNN, we use the $Lab$ color features of the $N=h{\times}w$ pixels, denoted as $F_{c}\in\mathbb{R}^{N{\times}3}$. The pixel coordinates are also given as input, but instead of the 2D pixel positions, we provide the 3D spherical coordinates $F_{s}\in[-1,1]^{N{\times}3}$. To match the coordinates domain, we normalize the Lab features $F_{c}$ to also lie in $[-1,1]$.
@@ 2.3.3 Horizontally circular clustering
For example, Fig. 3(a) shows the result obtained by using a standard zero padding strategy in the CNN layers. With $3{\times}3$ convolution kernels, the features extracted for pixels at $j=0$ and $j=w-1$ are not consistent with the ones of their neighborhood, which disrupts the selection of their closest superpixel among the $9$ closest. When computing the hard clustering association, border pixels are generally associated to a disconnected region resulting in the appearance of an artificial vertical border in the spherical space, as for planar methods.
@@
Figure 3: Impact of the circular padding on the superpixel segmentation. (a) With standard zero padding, the CNN features of pixels at $j=0$ and $j=w-1$ are not consistent with neighborhood, leading to a vertical border in the spherical space, as for planar methods. (b) With circular padding, the features remain consistent on the borders and the method is able to fully consider the geometry of the omnidirectional images.
@@ 2.3.5 Region connectivity
Finally, we use the panoramic stretch approach of [21] to introduce spatial distortions. To stretch a 360° image, $[x,y,z]$ coordinates are simply multiplied by a respective factor $[k_{x},k_{y},k_{z}]$ and projected back to the sphere. Pixel values are then computed using bilinear interpolation. Since setting $k_{z}$ would affect the projection of $x$ and $y$ values the same way, authors propose to only set $k_{x}$ and $k_{y}$ parameters. The 3D coordinates maps in Fig. 1 represent the image area that would be affected by increasing one of the parameters. For instance, setting $k_{y}<1$ would zoom on the region where $y$ values are close to -1 and 1 (see Fig. 4(right)). We refer the reader to [21] to more details on the stretching algorithm and to our supp. mat. for additional examples.
@@ 3.1.1 Datasets
{comment}
In [9], a synthetic omnidirectional road dataset from [sekkat2020omniscape] was considered along with the PSD. However, using these images to compare deep-learning based methods would lead to a significant bias since the image content diversity is low and they all feature a motorbike at the exact same position in each frame. To further demonstrate the performance of DSS, we choose to consider for the first time in spherical superpixel methods evaluation, the Wild PAnoramic Semantic Segmentation (WildPASS) dataset [28], containing 500 omnidirectional natural road images. We resize the images to $512{\times}1024$ and split the dataset into respectively 300, 100 and 100 images for train, validation and test sets.
@@ 3.1.3 Evaluation metrics
with $\epsilon$ a distance threshold set to $2$ pixels [10], and $\delta[a]=1$ when $a$ is true and 00 otherwise. Since it only measures recall, BR should be compared to the Contour Density (CD), i.e., the proportion of border pixels of the generated superpixels.
@@ 3.2 Ablation Study
{comment}
3.2.1 360 geometry
Si le temps de faire tourner le 3). On a le modèle de base. Ok c’est mieux en réentrainant mais peut-être pas parce que SPS peu d’images et on crop. On prend en compte la geometry 360 avec full image c’est mieux 3) On a jusqu’au bout en faisant du circular padding 4)
1 = 2 < 3 < 4
1) SSN trained on BSD
2) SSN trained on SPS
3) DSS trained on SPS w/o circular padding
4) DSS trained on SPS w/ circular padding
3.2.2 Data Augmentation
@@
{comment}
Table 2: Ablation study multiple hypothesis. Results are presented in AUC
Gaussian Blur + Noise Horizontal crop & mirror Panoramic strecth Circular padding ASA F CD/BR G-GR
- - - - XXX XXX XXX XXX
- - - XXX XXX XXX XXX
- - XXX XXX XXX XXX
- XXX XXX XXX XXX
XXX XXX XXX XXX
@@ 3.3.2 Evaluation of performance
{comment}
Refer to caption Refer to caption
(a) ASA on PSD images (b) ASA on noisy PSD images
Refer to caption Refer to caption
(a) ASA on PASS images (b) legend
Figure 6: A compléter.

### v2

@@ 1 Introduction
[…]
@@ 2.1 Superpixel Sampling Network
The SSN model takes as input images of size $N=h\times w$, represented with 5 channels corresponding to Lab color features (3 channels) and $xy$ pixel coordinates (2 channels). The goal of the model is to learn deep features that are more suitable to perform a differentiable clustering into superpixels. To achieve this, the SSN model uses a CNN composed of three blocks, each with two convolutional layers, batch normalization and ReLU activation, with a max pooling layer applied after each block. For the output, feature maps of each block are upsampled to the original image size (if necessary, for the second and third blocks) and concatenated. The original Lab and xy features are also concatenated into the output feature maps, resulting in $D$-dimensional pixel features (i.e., 5 channels from input features and D5D-5 learned deep features). In practice, the SSN model used $D=20$ in their experiments. For more details about this architecture, the reader can refer to [12].
@@ 2.3.1 Features and superpixels initialization
As input for the CNN, we use the $Lab$ color features of the $N=h{\times}w$ pixels, denoted as $F_{c}\in\mathbb{R}^{N{\times}3}$. The pixel coordinates are also given as input, but instead of the 2D pixel positions, we provide the 3D spherical coordinates $F_{s}\in[-1,1]^{N{\times}3}$. To match the coordinates domain, we normalize the Lab features $F_{c}$ to also lie in [1,1][-1,1].
@@ 2.3.3 Horizontally circular clustering
For example, Fig. 3(a) shows the result obtained by using a standard zero padding strategy in the CNN layers. With $3{\times}3$ convolution kernels, the features extracted for pixels at $j=0$ and j=w1j=w-1 are not consistent with the ones of their neighborhood, which disrupts the selection of their closest superpixel among the $9$ closest. When computing the hard clustering association, border pixels are generally associated to a disconnected region resulting in the appearance of an artificial vertical border in the spherical space, as for planar methods.
@@
Figure 3: Impact of the circular padding on the superpixel segmentation. (a) With standard zero padding, the CNN features of pixels at $j=0$ and j=w1j=w-1 are not consistent with neighborhood, leading to a vertical border in the spherical space, as for planar methods. (b) With circular padding, the features remain consistent on the borders and the method is able to fully consider the geometry of the omnidirectional images.
@@ 2.3.5 Region connectivity
Finally, we use the panoramic stretch approach of [21] to introduce spatial distortions. To stretch a 360° image, $[x,y,z]$ coordinates are simply multiplied by a respective factor $k_{x},k_{y},k_{z}]$ and projected back to the sphere. Pixel values are then computed using bilinear interpolation. Since setting $k_{z}$ would affect the projection of $x$ and $y$ values the same way, authors propose to only set $k_{x}$ and $k_{y}$ parameters. The 3D coordinates maps in Fig. 1 represent the image area that would be affected by increasing one of the parameters. For instance, setting $k_{y}<1$ would zoom on the region where $y$ values are close to -1 and 1 (see Fig. 4(right)). We refer the reader to [21] to more details on the stretching algorithm and to our supp. mat. for additional examples.
@@ 3.1.1 Datasets
To further demonstrate the performance of DSS, we choose to consider for the first time in spherical superpixel methods evaluation, the Wild PAnoramic Semantic Segmentation (WildPASS) dataset [28], containing 500 omnidirectional natural road images. We resize the images to $512{\times}1024$ and split the dataset into respectively 300, 100 and 100 images for train, validation and test sets.
@@ 3.1.3 Evaluation metrics
with $\epsilon$ a distance threshold set to $2$ pixels [10], and $\delta[a]=1$ when $a$ is true and 0 otherwise. Since it only measures recall, BR should be compared to the Contour Density (CD), i.e., the proportion of border pixels of the generated superpixels.